Could you please help me solve this problem?
Three drill bits are marked 3/8, 5/16, and 11/32. Which drill bit is larger?
Thanks,
Sher

Found 2 solutions by checkley71, Nate:
Answer by checkley71(8403)

TO COPMPARE FRACTION YOU NEED TO HAVE A COMMON DENOMINATOR THUS 3/8=12/32 & 5/16=10/32 & 11/32=11/32 THUS THE LARGEST NUMERATOR IS 12 THUS THE 3/8 DRILL IS THE LARGEST.

Answer by Nate(3500)

3/8, 5/16, and 11/32
3(4)/8(4), 5(2)/16(2), and 11/32
12/32, 10/32, and 11/32
(3/8) is the largest drill bit